OpenStudy (anonymous):

Find the value of the series (2^n+1)/3^n.

OpenStudy (anonymous):

From n=1 to infinity.

OpenStudy (anonymous):

@dumbcow @iambatman

OpenStudy (anonymous):

@myininaya @RadEn

OpenStudy (reemii):

is it A) \((2^n+1)/3^n\) or B) \(2^{n+1}/3^n\) ?

OpenStudy (anonymous):

B.

OpenStudy (reemii):

ok. When seeing these powers, you might thnk about the geometric formula \[\sum_{n=0}^\infty x^n = \frac{1}{1-x}\quad\text{if } 0<x<1. \] and try to make your series look like that (factor out a few things).

OpenStudy (anonymous):

So you mean that this is geometric series?

OpenStudy (dumbcow):

\[\frac{2^{n+1}}{3^n} = 2 (\frac{2}{3})^n\]

OpenStudy (anonymous):

Got it. Thank you so much, dumbcow.